Two moles of helium gas initially at 273 K and 0.31 atm are compressed isothermally to 1.55 atm.

1. Find the final volume of the gas. Assume that helium behaves as an ideal gas. The universal gas constant is 8.31451 J/K · mol.
Answer in units of m3.

2. Find the work done by the gas. Answer in units of kJ.

3. Find the thermal energy transferred. Answer in units of kJ.
1
Expert's answer
2019-04-08T09:22:18-0400

Ideal gas law:

P2*V2 = n*R*T

Solve for V2:

V2 = n*R*T/P2

 = 2*0.0831*273 /1.55

 =29.2 litre

=0.0291m^3

A) V2 = n*R*T/P2

B) W = n*R*T*ln(P1/P2)

C) Q = n*R*T*ln(P1/P2)

Data (pressures translated to kPa):

n:= 2 moles; R:=8.3145 J/mol-K; T:=273K; P1:=31.41075kPa; P2:157.0538kPa;

Unit note: J = kPa Liters...thus formula for V2 will return Liters.

B) W = -7.3kJ, negative indicates work done on the gas by an outside agent.

C) Q =-7.30kJ, negative indicates heat released to surroundings